The Jurassic Zoo charges ​$15 for each adult admission and ​$3 for each child. The total bill for the 75 people from a school trip was ​$477. How many adults and how many children went to the​ zoo?
Answer by addingup(3677)

a+c = 75 subtract c from both sides:
a = 75-c
and
15a+3c = 477
substitute for a:
15(75-c)+3c = 477
1125-15c+3c = 477
-12c = -648
c = 54 this is the number of children that went to the zoo.
And 75-54 = 21 adults
---------------------------
Check:
54*3 = 162
21*15 = 315
Total= 477 Correct